### v2

# Probing Hernquist dark matter with black hole shadows: A comprehensive study of various accretions

Abstract
The shadow of a black hole is critically dependent on the surrounding accreting matter. We investigate the observational signatures of a Schwarzschild black hole embedded in a Hernquist dark matter (DM) halo under three distinct accretion scenarios: a geometrically thin disk, a static spherical flow, and an infalling spherical flow. For the thin disk model, we find that direct emission dominates the total observed intensity, while the size and brightness of the lensing and photon rings serve as sensitive probes of the Hernquist DM parameters. Under spherical accretion, the Hernquist DM halo significantly enlarges the photon sphere. This results in an observable shadow that is approximately $2\%$ to $30\%$ larger than in the vacuum case, though this increase in size is accompanied by a considerable decrease in overall image brightness. Furthermore, the Doppler de-boosting effect in the infalling scenario produces a markedly darker image than its static counterpart. Our results demonstrate that the size and brightness profile of a black hole shadow provide a powerful observational tool to probe and constrain the distribution of dark matter in galactic centers.
Hernquist dark matter, black hole shadow, observational appearance.
I Introduction
The observation of black hole shadows has entered a new era, establishing them as a ground-breaking tool for researching powerful gravitational forces. The fundamental nature of gravity has been further illuminated by analysis of the critical behaviour of photons close to black holes. Recent advances in the Event Horizon Telescope’s (EHT) imaging of the supermassive black holes $\text{M87}^{*}$ Collaboration (2019a, b, c, d, e, f, 2021a, 2021b, 2023) and $\text{SgrA}^{*}$ Ghez et al. (2005); Boehle et al. (2016); Akiyama et al. (2022a, b, c, d, e, f, 2024a, 2024b) have shown shadow features that not only support general relativity (GR) Hawking (1972); Bardeen et al. (1973a); Wald (1994, 2010) predictions but also provide a novel window of probing black holes. The black hole shadow, also known as the photon capture zone, is the centre dark area shown in these pictures Perlick and Tsupko (2022). This is encircles a compact, asymmetrical and dazzling circular structure. This phenomenon, known as the photon ring, results from light beams from infinity being deflected by the strong gravitational field as they approach the black hole. In backward ray-tracing methods, light rays that approach the critical impact parameter asymptotically converge to the bound photon orbit. The so-called photon sphere, which projects the shadow of the black hole, is made up of this confined photon orbit. An angular radius equation for the shadow of Schwarzschild black holes has been established by Synge and Luminet using a crucial impact parameter of $b_{p}=3\sqrt{3}M$ Synge (1966); Luminet (1979). The first image of a black hole surrounded by a thin accretion disk was analytically produced by the authors in Ref.Luminet (1979), displaying primary and secondary images that emerge beyond the black hole’s shadow. Bardeen later studied the D-shaped shadow of Kerr black holes, showing that the shape of the shadow can be changed by the spin of the black hole Bardeen et al. (1973b); Chandrasekhar (1998). Additionally, a great deal of research has been done on black hole shadows in higher-dimensional spacetime and other modified gravity theories Amarilla et al. (2010); Amarilla and Eiroa (2012); Amir et al. (2018); Mizuno et al. (2018); Eiroa and Sendra (2018); Banerjee et al. (2020); Bambi et al. (2019); Vagnozzi et al. (2023); Khodadi et al. (2024). In contrast to single shadows, the research has also suggested multiple shadows for wormholes and black holes Huang et al. (2016); Guerrero et al. (2021). While the finding of the EHT mainly relates to black holes in GR, it provides a great deal of opportunity to investigate other compact objects in theories other than GR.
To investigate the main characteristics of black hole pictures and find possible indications of novel physics, it is helpful to set up a basic accretion model. Schwarzschild black holes were studied for the spherically symmetric accretion model, another accretion scenario, and it was shown that under such accret ion settings the shadow has resilient properties where its shape and size are determined by spacetime instead of accretion details Gralla et al. (2019); Tsukamoto et al. (2014); Tsukamoto (2018). Further investigations of optically thin and geometrically thin accretion models show that the size of the shadow is strongly related to the position of the accretion Zeng and Zhang (2020); Zeng et al. (2020); Li and He (2021a); He et al. (2022); Peng et al. (2021); Zeng et al. (2022, 2025). Direct emission, lensing rings, and photon rings are always present in the bright areas outside the shadow. These characteristics make it possible to distinguish GR black holes from other compact objects or black holes in modified gravity theories through observation Zeng and Zhang (2020); Zeng et al. (2020); Li and He (2021a); He et al. (2022); Peng et al. (2021); Zeng et al. (2022, 2025). Consequently, EHT observations of black hole shadows are extremely important for constraining parameters in modified gravity theories and testing fundamental physics Shaikh et al. (2019); Davoudiasl and Denton (2019); Roy and Yajnik (2020).
The environment surrounding the black hole, in particular the distribution of DM, offers an additional convincing explanation for possible departures from conventional predictions, in addition to modified gravity theories. The discovery of huge elliptical and spiral galaxies was the first significant development in the seek for DM Rubin et al. (1980). Persic found that about $90\%$ of a galaxy’s mass is made up of DM Persic et al. (1996). Estimating DM’s contribution close to the galactic centre is especially important since there is strong evidence that DM halos around astrophysical black holes Collaboration (2019a, f); Sofue (2013); Boshkayev and Malafarina (2019); Jusufi et al. (2019); Konoplya (2019); Konoplya and Zhidenko (2022); Hou et al. (2018a); Liang et al. (2023); Gohain et al. (2024); Övgün and Pantig (2025); Jha (2025). Several DM distributions are covered by the Dehnen density distribution, which is frequently used to describe dwarf galaxies Dehnen (1993); Mo et al. (2010). These investigations usually use the $\text{Dehnen}-(1,4,1)$ type, which is the Hernquist DM distribution. The Hernquist decay is more consistent with the dynamical evidence of dwarf elliptical galaxies under observational restrictions. Nevertheless, there is still a dearth of comprehensive investigation into how Hernquist DM halos alter the Schwarzschild black hole shadow’s observable properties. Although the effects of Hernquist DM halos on the photon sphere radius and shadow radius of Schwarzschild black holes are discussed in Ref.Jha (2025); Ahmed et al. (2025); Feng and Zhang (2025), accretion models are not taken into account. Furthermore, the direct contribution of important factors like centre density and core radius to lensing rings and photon rings is not quantified in accretion model studies, which mostly concentrate on vacuum or uniform DM backgrounds Gralla et al. (2019); Zeng and Zhang (2020); Zeng et al. (2020); Li and He (2021a); He et al. (2022); Peng et al. (2021); Zeng et al. (2022, 2025); Shaikh et al. (2019); Shi and Cheng (2024); Feng and Zhang (2025); Amir et al. (2018); Eiroa and Sendra (2018); Amarilla and Eiroa (2012); Konoplya (2019); Mizuno et al. (2018); Meng et al. (2023). There is currently no systematic investigation that combines comprehensive accretion models with a Hernquist DM halo. We cannot use black hole shadows as accurate probes to restrict DM characteristics due to this gap.
In this work, we present the first comprehensive characterization of the shadow of Schwarzschild-Hernquist black holes with different accretion models. In order to clarify the associated impacts of Hernquist DM halo parameters on the photon sphere and observed brightness, we thoroughly investigated three thin disk accretion models and two spherical accretion models. Apparently, different accretion regimes produce different shadow pictures under the same conditions. In particular, pictures made up of the shadow, lensing ring, and photon ring are created via thin disk accretion. The shadow and photon ring are spherically symmetric for spherical accretion. The observed intensity is affected differently by static and falling accretion, which significantly changes the brightness and radius of the shadow. A darker shadow than under static conditions results from the extra Doppler effect introduced by falling accretion Li and He (2021b, a); Zeng and Zhang (2020); Zeng et al. (2020, 2022, 2025). We close a gap in the references on Hernquist metrics and accretion models by revealing scaling rules for shadow size and brightness as factors of the Hernquist DM halo distribution. This offers an observable diagnostic standard for differentiating amongst Hernquist DM characteristics.
The structure of this paper is as follows: Sec.II derives the zero-geodesic equation under the Hernquist metric, defining the photon sphere radius and collision parameter; Sec.III analyses the thin disk accretion model, including the contribution regions of direct radiation, lensing rings, and photon rings, and relates the emission intensity to the observed appearance; Sec.IV examines static and falling spherical accretion models, contrasting how varying Hernquist DM halo parameters affect shadow brightness; finally, Sec.V synthesises the findings, emphasising the potential interpretative power of Hernquist DM halos for shadows and outlining their prospects for detecting DM distributions in galactic centres.
II Light deflection and photon orbit
In this work, we model a Schwarzschild black hole immersed in a DM halo that is described by the Hernquist density profile Dehnen (1993); Mo et al. (2010),
$\rho(r)=\rho_{c}\left(\dfrac{r}{r_{s}}\right)^{-1}\left[1+\dfrac{r}{r_{s}}\right]^{-3},$ (1)
where $\rho_{c}$ and $r_{s}$ represent the center density and core radius of the DM halo, respectively. At short radii $(r\ll r_{s})$, this distribution shows an isothermal feature of $\rho\propto r^{-1}$, but at high radii $(r\gg r_{s})$, it decays to $\rho\propto r^{-4}$Jha (2025). For details of the computational process, we refer the reader to Refs.Mo et al. (2010); Yang et al. (2024); Xu et al. (2018). Here, the metric of a Schwarzschild black hole immersed in the Hernquist DM halo surroundings can be written as follows Jha (2025); Ahmed et al. (2025); Feng and Zhang (2025):
$\mathrm{d}s^{2}=-f(r)\mathrm{d}t^{2}+\dfrac{\mathrm{d}r^{2}}{f(r)}+r^{2}\left(\mathrm{d}\theta^{2}+\sin^{2}\theta\mathrm{d}\varphi^{2}\right),$ (2)
where
$f(r)=1-\dfrac{2M}{r}-\dfrac{4\pi\rho_{c}r_{s}^{3}}{r+r_{s}}.$ (3)
Here, the Newton constant $G$ has been set to $1$ for simplicity. Naturally, this black hole metric (3) reduces to the Schwarzschild black hole in the limit $\rho_{c}\to 0$ or $r_{s}\to 0$. Our goal is to investigate the Schwarzschild-Hernquist black hole’s null geodesic. We restrict our analysis to the equatorial plane, $\theta=\pi/2$ and $\dot{\theta}=0$Synge (1966); Bardeen et al. (1972); Gralla and Lupsasca (2020). Therefore, the Lagrangian $\mathcal{L}$ for a particle in this spacetime can be easily obtained,
$\mathcal{L}$ $=\frac{1}{2}g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}$ (4)
$=\frac{1}{2}\left[-f(r)\dot{t}^{2}+\dfrac{\dot{r}^{2}}{f(r)}+r^{2}\dot{\varphi}^{2}\right],$ (5)
where, $\dot{x}^{\mu}=\partial x^{\mu}/\partial\chi$ is the four-velocity of the photon and $\chi$ is the affine parameter. Since the coefficient of the metric equation cannot be directly found using the $t$ and $\theta$ coordinates, energy and angular momentum are represented by the conserved variables $E$ and $L$, respectively. That’s Shaikh et al. (2019); He et al. (2022); Li and He (2021a); Shi and Cheng (2024); Feng and Zhang (2025)
$-E$ $=\dfrac{\partial\mathcal{L}}{\partial\dot{t}}=-f(r)\dot{t},$ (6)
$L$ $=\dfrac{\partial\mathcal{L}}{\partial\dot{\varphi}}=r^{2}\dot{\varphi}.$ (7)
By taking into account that $g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}=0$ for the null geodesic, and then using Eqs.(6)-(7) to solve Eq.(4), it produces Amir et al. (2018); Eiroa and Sendra (2018); Shaikh et al. (2019); He et al. (2022); Li and He (2021a)
$\dot{t}$ $=\dfrac{1}{b}\left(1-\dfrac{2M}{r}-\dfrac{4\pi\rho_{c}r_{s}^{3}}{r+r_{s}}\right)^{-1},$ (8)
$\dot{\varphi}$ $=\pm\dfrac{1}{r^{2}},$ (9)
$\dot{r}^{2}$ $=\dfrac{1}{b^{2}}-\dfrac{1}{r^{2}}\left(1-\dfrac{2M}{r}-\dfrac{4\pi\rho_{c}r_{s}^{3}}{r+r_{s}}\right).$ (10)
In this case, the impact parameter is $b=|L|/E$, and we redefine the affine parameter $\chi\to\chi/|L|$, and the sign $\pm$ denotes the light ray’s clockwise and anticlockwise directions. When the effective potential $V_{\text{ph}}$ is included, we may rewrite Eq.(10) as
$\dot{r}+V_{\text{ph}}=\dfrac{1}{b^{2}},$ (11)
where
$V_{\text{ph}}=\dfrac{1}{r^{2}}\left(1-\dfrac{2M}{r}-\dfrac{4\pi\rho_{c}r_{s}^{3}}{r+r_{s}}\right).$ (12)
The conditions for a photon sphere orbit are $\dot{r}=0$ and $\ddot{r}=0$, which translate to
$V_{\text{ph}}$ $=\dfrac{1}{b^{2}},$ (13)
$\dfrac{\mathrm{d}V_{\text{ph}}}{\mathrm{d}r}$ $=0.$ (14)
This equation shows that when the photon sphere radius $r_{p}$ changes, the black hole’s effective potential reaches a maximum, which corresponds to real and positive roots. Consequently, the connection between the critical impact parameter $b_{p}$ and the photon sphere radius $r_{p}$ is as followsGralla and Lupsasca (2020); Li and He (2021b); Fathi and Cruz (2023); He et al. (2022); Shi and Cheng (2024):
$r_{p}^{2}$ $=b_{p}^{2}f(r),$ (15)
$\dfrac{r_{p}^{3}}{2}\dfrac{\mathrm{d}f(r)}{\mathrm{d}r}$ $=b_{p}^{2}f(r)^{2}.$ (16)
The radius $r_{p}$ and the impact parameter $b_{p}$ of the photon sphere are represented in Eqs.(15)-(16), respectively. Tab.1 displays numerical results for the event horizon radius $r_{h}$, photon sphere radius $r_{p}$, and impact parameter $b_{p}$ by choosing various values for $\rho_{c}$ and $r_{s}$. It is evident that as $\rho_{c}$ and $r_{s}$ increase, the event horizon radius $r_{h}$, the photon sphere radius $r_{p}$, and the impact parameter $b_{p}$ all rise, exceeding their values in the Schwarzschild case. This increase is a direct physical consequence of the additional mass-energy contributed by the DM halo, which enhances the spacetime curvature outside the event horizon. To maintain an unstable circular orbit within this stronger gravitational field, a photon must possess a larger angular momentum. For a distant observer, this corresponds to a larger critical impact parameter $b_{p}$, and the radius of the orbit itself $r_{p}$ is consequently shifted outwards. In essence, the DM halo enlarges the ability of the black hole for capturing photons.
Fig.1 shows the effective potential $V_{\text{ph}}$ under various $\rho_{c}$ and $r_{s}$ conditions. As an example, for case $\rho_{c}M^{2}=0.4/M^{2}$ and $r_{s}/M=0.3M$ , Fig.1 shows that the effective potential $V_{\text{ph}}$ is zero near the event horizon and only occurs when $r\geq r_{p}$. It grows monotonically from this point and reaches its maximum at the photon sphere $r_{p}$. After then, $V_{\text{ph}}$ progressively decreases to 0 as $r$ rises from the photon sphere towards infinity. The effective potential $V_{\text{ph}}$ naturally forms a potential barrier for incident light rays. Some light rays traveling inward from infinity are reflected by this barrier. The trajectories of these rays correspond to those in Region 1 $(b>b_{p})$ in Fig.1. In Region 3 $(b<b_{p})$ of Fig.1, certain rays will not pass over the barrier and will eventually fall into the black hole. Additionally, in Region 2 $(b=b_{p})$, rays will asymptotically reach the photon sphere’s orbit before continuing indefinitely around the black hole.
Table 1: The numerical results of the event horizon $r_{h}$, the photon sphere $r_{p}$ and the impact parameter $b_{p}$ of Schwarzschild-Hernquist black holes with $M=1$.
$r_{h}/M$ $r_{p}/M$ $b_{p}/M$ $r_{h}/M$ $r_{p}/M$ $b_{p}/M$ $r_{h}/M$ $r_{p}/M$ $b_{p}/M$
$r_{s}/M=0.2$ $r_{s}/M=0.3$ $r_{s}/M=0.4$
$\rho_{c}M^{2}=0.4$ 2.03662 3.05545 5.29421 2.11888 3.18068 5.51839 2.27357 3.41689 5.94454
$\rho_{c}M^{2}=0.6$ 2.05497 3.08324 5.34332 2.17894 3.27188 5.68068 2.41395 3.63042 6.32576
$\rho_{c}M^{2}=0.8$ 2.07335 3.11107 5.39249 2.23937 3.36361 5.84369 2.55634 3.84675 6.71082
Figure 1: The effective potential $V_{\text{ph}}$ and the impact parameter $b_{p}$ of the Schwarzschild-Hernquist black holes.
In order to determine the trajectory of light rays, the equation of motion for photons is reformulated using Eqs.(9)-(10),
$\dfrac{\mathrm{d}r}{\mathrm{d}\varphi}=\pm r^{2}\sqrt{\dfrac{1}{b^{2}}-\dfrac{1}{r^{2}}\left(1-\dfrac{2M}{r}-\dfrac{4\pi\rho_{c}r_{s}^{3}}{r+r_{s}}\right)}.$ (17)
By substituting $u\equiv 1/r$, the equation above may be written as
$\frac{\mathrm{d}u}{\mathrm{d}\varphi}=\sqrt{\frac{1}{b^{2}}-u^{2}\left[1-2Mu-\dfrac{4\pi\rho_{c}r_{s}^{3}u}{1+r_{s}u}\right]}.$ (18)
Via solving Eq.(18) numerically using a ray tracing code, Fig.2 shows the trajectories of light ray. The black hole is shown in Fig.2 as a solid disk. Photons are caught by the black hole when the collision parameter $b<b_{p}$ (Region 3 in Fig.1). The blue lines in Fig.2 represent their paths. On the other hand, photons are redirected and moved towards infinity if $b>b_{p}$ (Region 1 in Fig.1), with their paths matching the yellow lines in Fig.2. Photons circle the black hole indefinitely when $b=b_{p}$ (Region 2 in Fig.1). The red lines in Fig.2 represent their respective paths. It is clear that the red lines eventually form a circle, and the location of this circle exactly matches the black hole’s photon sphere orbit. In Fig.2, all incoming light rays are initialized to be parallel to the horizontal axis, which is a valid choice since the Schwarzschild-Hernquist spacetime is asymptotically flat. Moreover, it is impossible to ignore the substantial impact that the center density $\rho_{c}$ and the core radius $r_{s}$ of the Hernquist DM halo have on the photon paths.
Figure 2: The behaviour of light trajectories under various $\rho_{c}$ and $r_{s}$ in the polar coordinate system $(r,\varphi)$.
III Observational features of thin disk emission
In astrophysical settings, black holes are typically surrounded by large amounts of accreting matter. This makes the study of shadows cast by Schwarzschild-Hernquist black holes particularly relevant from an observational perspective. In this section we will investigate the shadows and rings of a Schwarzschild-Hernquist black hole surrounded by a disk-like accretion disk that is optically and geometrically thin.
III.1 Direct emission, lensing ring and photon ring
According to Ref.Gralla et al. (2019), radiation with different impact parameters $b$ would cause the black hole to look differently to a distant observer when tracing rays backwards from an observer towards the black hole. We may use ray-tracing procedure to determine the optical appearance of a Schwarzschild-Hernquist black hole, by calculating the total deflection in azimuthal angle $\varphi$, for a photon traveling from the source to the observer. The number of orbits is then given by Gralla et al. (2019); Gralla and Lupsasca (2020)
$n(b)=\dfrac{\varphi}{2\pi},$ (19)
Eq.(19) depends on the impact parameter $b$. The number of orbits $n(b)$ is determined by how close the impact parameter b is to the critical value $b_{p}$. It also depends on the metric (3) of the Schwarzschild-Hernquist black hole.
Figure 3: The number of orbits $n$ vs the impact parameter $b$ is shown for various Hernquist DM parameters $\rho_{c}$ and $r_{s}$.
Figure 4: Photon behaviour as a function of impact parameter $b$ in the Schwarzschild-Hernquist black hole.
Table 2: The region of direct emission, lensing ring and photon ring is related to the impact parameters $b$ for different Hernquist DM parameters with $M=1$.
Type $r_{s}/M=0.2$ $r_{s}/M=0.4$
$\rho_{c}M^{2}=0.4$
Direct emission $b<5.10927$ or $b>6.28594$ $b<5.73082$ or $b>7.08226$
Lensing ring
$5.10927<b<5.28566$
or $5.32673<b<6.28594$
$5.73082<b<5.93441$
or $5.98264<b<7.08226$
Photon ring $5.28566<b<5.32673$ $5.93441<b<5.98264$
$\rho_{c}M^{2}=0.8$
Direct emission $b<5.20362$ or $b>6.40451$ $b<6.46474$ or $b>8.01367$
Lensing ring
$5.20362<b<5.38374$
or $5.42573<b<6.40451$
$6.46474<b<6.69895$
or $6.75509<b<8.01367$
Photon ring $5.38374<b<5.42573$ $6.69895<b<6.75509$
In favour of the number of orbits $n(b)$, the emission can be categorized into three distinct zones Gralla et al. (2019); Gralla and Lupsasca (2020); Zeng and Zhang (2020); Zeng et al. (2020, 2022, 2025); Li and He (2021b, a). These are lensing rings, photon rings, and direct emission rings corresponding to $n<3/4$, $3/4<n<5/4$, and $n>5/4$, respectively. These zones correspond to light rays that intersect the equatorial plane once, twice, or more than twice, respectively Gralla et al. (2019); Gralla and Lupsasca (2020). For various values of the parameters $\rho_{c}$ and $r_{s}$, Fig.3 shows the functional dependence of $n(b)$ on the impact parameter $b$. The domains of direct emissions, lensing rings, and photon rings are shown by different colors. We have indicated the ring and the direct emission locations in Fig.4. The function n(b)diverges as bapproaches the critical value $b_{p}$, indicating the existence of unstable photon orbits, as shown in Fig.3. The curves for larger DM parameters $\rho_{c}$ and $r_{s}$ are shifted to the right, consistent with the increase in $b_{p}$ reported in Tab.1. Tab.2 lists the ranges of the impact parameter bcorresponding to direct emission, the lensing ring, and the photon ring for various values of $\rho_{c}$ and $r_{s}$. It is evident from the data that the range of the impact parameter $b$ for all emission types increases as the parameters increase. Furthermore, the total number of photon orbits $n(b)$ will peak inside a small area when the impact parameter $b$ gets closer to the critical condition $b\sim b_{p}$. The extremely narrow ranges for the photon ring listed in Tab.2 correspond to the thin red band visible in Fig.4. The angular size of the shadow also increases with larger parameters. However, this enlargement does not directly translate to an increase in the brightness of the ring; the observed intensity depends on the complex interplay between the emission profile and the light bending.
III.2 Transfer functions
We now study the emission from a thin disk viewed in a face-on orientation. In this case, we suppose that the thin disk is located on the Schwarzschild-Hernquist black hole’s equatorial plane and that its emission is isotropic in the rest frame of the emitting material. The specific intensity measured by a stationary observer at infinity, neglecting absorption and scattering, is given by Zeng and Zhang (2020); Zeng et al. (2020, 2022, 2025); Li and He (2021b, a)
$I(r)=\left(1-\dfrac{2M}{r}-\dfrac{4\pi\rho_{c}r_{s}^{3}}{r+r_{s}}\right)^{3/2}I_{\text{e}}(r).$ (20)
The total observed intensity, which integrated over all frequencies, is obtained by integrating Meng et al. (2023)
$I_{\text{o}}\left(r\right)$ $=\int I(r)\mathrm{d}\nu$ (21)
$=\left(1-\dfrac{2M}{r}-\dfrac{4\pi\rho_{c}r_{s}^{3}}{r+r_{s}}\right)^{2}I_{\text{e}}(r),$
where the total emission intensity is given by
$I_{\text{emi}}(r)=\int I_{\text{e}}(r)\mathrm{d}\nu_{\text{e}}.$ (22)
Since a light ray may intersect the accretion disk multiple times, each intersection contributes to the observed intensity. Therefore, the total observed intensity is a sum over these intersections:
$I_{\text{obs}}\left(r\right)=\left(1-\dfrac{2M}{r}-\dfrac{4\pi\rho_{c}r_{s}^{3}}{r+r_{s}}\right)^{2}\sum_{n}I_{\text{emi}}(r)|_{r=r_{n}(b)}.$ (23)
Figure 5: The first three transfer functions for the Schwarzschild-Hernquist black hole.
The transfer function $r_{n}(b)$ encodes the radial coordinate on the disk of the $n$-th intersection of a light ray with impact parameter $b$. Furthermore, the transfer function’s slope, often called the demagnification factor, is $\mathrm{d}r/\mathrm{d}b$, and it determines the magnification of the image of the disk Gralla et al. (2019); Li and He (2021b); Fathi and Cruz (2023); He et al. (2022); Shi and Cheng (2024); Zeng and Zhang (2020). In Fig.5, the transfer function for direct emission $(n=1)$ is shown by the blue line. Since the direct imaging profile is the redshifted source profile, the slope $\mathrm{d}r/\mathrm{d}b$ of the blue line remains approximately constant and close to 1 despite changes in the Hernquist DM parameters. For the lensing ring $(n=2)$, the transfer function is shown by the yellow line. The slope of the second transfer function rises to a very high value when the impact parameter b gets closer to the critical condition $b\sim b_{p}$. For the photon ring $(n=3)$, the red line represents the transfer function. The slope of the red line approaches infinity, indicating extreme demagnification of the disk’s image. This implies that direct emission dominates the total observed flux, with the photon ring and lensing ring contributing only a small fraction. Notably, the contribution of transfer functions with $n\geq 4$ has been discovered to provide some interferometric featuresJohnson et al. (2020), but it is insignificant and negligible. Therefore, we consider only the first three transfer functions $n=1,2,3$.
III.3 Observational features
To compute the observational appearance of the thin disk, we employ three phenomenological emissivity models Wang et al. (2022); Yang et al. (2023); Li and He (2021a); Zeng and Zhang (2020); Zeng et al. (2020, 2022, 2025). It is important to note that these models are simplified representations and are strategically chosen to represent a diverse range of plausible astrophysical scenarios from efficient, geometrically thin disks to inefficient, thick flows Meng et al. (2023); Gan et al. (2021); Guo et al. (2022); Chen et al. (2022). This approach allows us to systematically explore how the observational signatures of the Hernquist DM halo depend on the spatial distribution of the emitting matter. While the precise quantitative features of the image, such as peak brightness, are model-dependent, we expect the qualitative findings, particularly the scaling of the shadow and photon ring diameters with DM parameters, to be robust across different emission profiles.
Model 1 is designed to represent a standard, geometrically thin, optically thick accretion disk, analogous to the classical Novikov-Thorne model. The emissivity is sharply peaked at the innermost stable circular orbit (ISCO), which is the radius where gravitational energy is most efficiently released, and follows a rapid power-law decay outwards. This profile is characteristic of radiatively efficient accretion scenarios, often seen in luminous active galactic nuclei (AGN).
$I_{\text{emi}}(r)=\left\{\begin{array}[]{ll}\left[\dfrac{1}{r-\left(r_{\text{ISCO}}-1\right)}\right]^{2},&r>r_{\text{ISCO}}\\ 0,&r\leq r_{\text{ISCO}}\end{array}\right.$. (26)
Model 2 serves as an exploratory case to probe the observational appearance when the emission peak is located extremely close to the unstable photon orbits at $r_{p}$. While not a standard astrophysical scenario, this profile could hypothetically arise from localized energy release, such as magnetic reconnection events occurring near the photon sphere. Its primary purpose is to test the most extreme gravitational lensing effects on an emission source situated at the boundary of photon capture.
$I_{\text{emi}}(r)=\left\{\begin{array}[]{ll}\left[\dfrac{1}{r-\left(r_{p}-1\right)}\right]^{3},&r>r_{p}\\ 0,&r\leq r_{p}\end{array}\right.$. (29)
Model 3 is chosen to represent a physically distinct scenario: a geometrically thick, optically thin, and radiatively inefficient accretion flow, such as an Advection-Dominated Accretion Flow (ADAF). This type of flow is believed to be present in low-luminosity AGN like $\text{M87}^{*}$ and $\text{SgrA}^{*}$. The emissivity profile is consequently broader, with a more gradual decay, reflecting emission that is spread over a much wider radial range rather than being concentrated at the inner edge.
$I_{\text{emi}}(r)=\left\{\begin{array}[]{ll}\dfrac{\frac{\pi}{2}-\tan^{-1}\left[r-\left(r_{\text{ISCO}}-1\right)\right]}{\frac{\pi}{2}+\tan^{-1}(r_{p})},&r>r_{h}\\ 0,&r\leq r_{h}\end{array}\right.$. (32)
We demonstrate the results for two representative parameter sets: $\rho_{c}M^{2}=0.4,r_{s}/M=0.2$ (Fig.6) and $\rho_{c}M^{2}=0.8,r_{s}/M=0.4$ (Fig.7). The emission profiles of the three models are shown in the left column of Figs.6 and 7 from top to bottom as Model 1, Model 2 and Model 3. The observed specific intensity, which is dependent on the impact parameter $b$, is given in the centre column. Additionally, the right column provides the two-dimensional observed image.
Figure 6: Observational view of the thin disk surrounding the Schwarzschild-Hernquist black hole with various emission profiles in case of $\rho_{c}M^{2}=0.4$ and $r_{s}/M=0.2$.
The first row of the left column in Figs.6 and 7 shows that the emissivity peaks near $r_{\text{ISCO}}$ and decays rapidly to zero with increasing $r$. In this instance, the photon sphere lies inside the emission zone. At the same time, two distinct peaks in the observed intensity profile, corresponding to the lensing ring and photon ring, can be identified, each of which can be separately differentiated due to the gravitational lensing effect during observation. However, compared to the direct emission peak, the photon sphere and the lensing ring have smaller observational areas and lower observational intensity peaks. Consequently, the total observed flux is dominated by direct emission, whereas the photon sphere contributes negligibly to the total flux and lensing ring emission contributes only a small fraction. The right-hand column displays the primary optical appearance, which is caused by direct emission, and the photon ring is barely visible.
Figure 7: Observational view of the thin disk surrounding the Schwarzschild-Hernquist black hole with various emission profiles in case of $\rho_{c}M^{2}=0.8$ and $r_{s}/M=0.4$.
The emission peak in the left column of the second row in Figs.6 and 7 is shifted outward. According to the results in the centre column, the measured intensity will rise to the first peak due to direct emission and then show a trend of progressive decrease as $r$ rises. Consequently, the observed intensity reaches a higher peak due to the combined emission from these rings. The measured intensity will therefore reach a new higher peak as a result of the combined effects of the photon ring, lensing ring, and direct emission. However, the lensing ring and photon ring are still limited to a small area. Nevertheless, the observation is still dominated by direct emission. The lensing ring still contributes only a small amount to the total flux, as can be seen from the right-hand column, although direct emission continues to be the dominant source.
As can be seen in the left column of the third row in Figs.6 and 7, when using Model 3, the emission extends inward, close to the event horizon $r_{h}$. Over a wider region, the lensing ring, photon ring, and direct emission overlap. Starting from just outside the event horizon, the observed intensity increases gradually, then rises sharply through the lensing ring region, and peaks within the photon ring region. The contribution from the lensing ring emission then causes the measured intensity to reach a lower peak. The observed intensity then starts to gradually decrease. Compared to the previous two models, the lensing ring’s contribution to the total observed intensity is more noticeable in this case. Additionally, a brilliant ring created by the combined effects of the lensing ring, photon ring, and direct emission is visible in the optical appearance.
The parameters of the Hernquist DM halo have a direct and significant impact on the observational features, as can be seen by a comparison of Figs.6 and 7. The photon and lensing ring radii both enlarge with increasing parameter, while their contribution to the total flux diminishes. Furthermore, for Model 3, the diameter of the bright ring in the image appears to increase slightly with increasing DM parameters. As the Hernquist DM halo parameter grows, the maximum measured intensity increases for all three emission scenarios ($2.12\%$ for Model 1, $6.22\%$ for Model 2 and $5.31\%$ for Model 3).
IV Shadows with spherical accretions
Having analyzed thin disk accretion, we now turn our attention to spherical accretion flows, which provide alternative astrophysical scenarios. We will investigate both static and infalling accretion models to determine how the Hernquist DM halo impacts the corresponding observational signatures.
IV.1 The static spherical accretion
We first study the shadow and photon ring in the scenario of a static spherical accretion surrounding a Schwarzschild-Hernquist black hole. The specific intensity observed at infinity is given by Jaroszynski and Kurpiewski (1997); Bambi (2013); Fathi and Cruz (2023)
$I_{\text{obs}}=\int_{\gamma}\mathcal{R}^{3}\mathcal{J}\left(\nu_{\text{emi}}\right)\mathrm{d}\ell_{\text{prop}},$ (33)
with
$\mathcal{J}\left(\nu_{\text{emi}}\right)\propto\dfrac{\delta\left(\nu_{\text{emi}}-\nu_{\text{fix}}\right)}{r^{2}},$ (34)
over the trajectory of emitted photons $\gamma$. Here, $\mathrm{d}\ell_{\text{prop}}$ is the infinitesimal proper length, $\mathcal{J}\left(\nu_{\text{emi}}\right)$ is the emissivity per unit volume in the rest frame of the accreting material, $\nu_{\text{emi}}$ is the emitted photon frequency, and $\nu_{\text{fix}}$ is the fixed, monochromatic emission frequency. In the Schwarzschild-Hernquist spacetime, the redshift factor is
$\mathcal{R}=\dfrac{\nu_{\text{obs}}}{\nu_{\text{emi}}}=\sqrt{f(r)},$ (35)
and the infinitesimal proper length is
$\mathrm{d}\ell_{\text{prop}}=\sqrt{\dfrac{1}{f(r)}+r^{2}\left(\dfrac{\mathrm{d}\varphi}{\mathrm{d}r}\right)^{2}}\mathrm{d}r.$ (36)
Combining Eqs.(33)-(36), the observed intensity for a static observer at infinity can be expressed as
$I_{\text{obs}}=\int_{\gamma}\dfrac{f(r)^{3/2}}{r^{2}}\sqrt{\dfrac{1}{f(r)}+r^{2}\left(\dfrac{\mathrm{d}\varphi}{\mathrm{d}r}\right)^{2}}\mathrm{d}r.$ (37)
The integrand in Eq.(37) represents the intensity contribution as a function of impact parameters $b$, which is plotted in Fig.8. Fig.9 displays the shadow and photon sphere of the Schwarzschild-Hernquist black hole by plotting the measured brightness versus $b$.
Figure 8: The specific intensity $I_{\text{obs}}$ of a static spherical accretion as observed by a distant observer.
Figure 9: The images of Schwarzschild-Hernquist black holes’ shaodow with static spherical accretion as observed by a distant observer.
The resulting observed intensity $I_{\text{obs}}(b)$ first increases with the impact parameter $b$, peaks near the critical value $b\sim b_{p}$, and then gradually decreases for larger $b$. The observational intensity asymptotically approaches zero as $b$ approaches infinity. Furthermore, we find that the measured intensity is strongly influenced by changes in the Hernquist DM halo. As seen in Fig.9, the observed intensity’s peak value decreases as $\rho_{c}$ and $r_{s}$ are increased. As shown in the images of Fig.9, the overall appearance becomes darker for larger values of $\rho_{c}$ and $r_{s}$. The image for the case $\rho_{c}M^{2}=0.4$ and $r_{s}/M=0.2$ is significantly brighter than the image for $\rho_{c}M^{2}=0.8$ and $r_{s}/M=0.4$. Interestingly, although the image becomes darker, the photon ring, which corresponds to the peak in $I_{\text{obs}}(b)$, expands to larger bvalues as $\rho_{c}$ and $r_{s}$ increase, consistent with the increase in $b_{p}$ reported in Tab.1. Keep in mind that the central intensity is not zero, but it is extremely modest, close to $b\sim 0$. Due to the small but non-zero probability of radiation originating from very close to the black hole to escape, the region inside the photon ring in the images of Fig.9 is not completely dark but exhibits a faint glow, brightest near the location of the photon ring itself.
IV.2 The infalling spherical accretion
Figure 10: The specific intensity $I_{\text{obs}}$ of a infalling spherical accretion as observed by a distant observer.
We next investigate the case of infalling spherical accretion around Schwarzschild-Hernquist black holes. The general formula for the observed intensity, Eq.(33), remains valid for infalling accretion. However, the redshift factor $\mathcal{R}$ must be modified to account for the velocity of the infalling matter. For infalling accretion, the redshift factor is given by Li and He (2021b, a); Fathi and Cruz (2023); Bambi (2013)
$\mathcal{R}=\dfrac{\kappa_{\mu}u^{\mu}_{\text{obs}}}{\kappa_{\nu}u^{\nu}_{\text{emi}}},$ (38)
The four-velocities for a distant static observer and for the infalling accreting matter are
$\bm{u}_{\text{obs}}=(1,0,0,0)$ (39)
and
$\bm{u}_{\text{emi}}=\left(\dfrac{1}{f(r)},-\sqrt{1-f(r)},0,0\right),$ (40)
respectively. In Eq.(38), the four-velocities of photons $\bm{\kappa}$ released from the accretion disk is represented by the co-vector, which has the same definition as in Eqs.(8)-(10). For purely radial accretion, we need to compute the ratio $\kappa_{r}/\kappa_{t}$. Using the photon’s equations of motion, this ratio is found to be Bambi (2013)
$\dfrac{\kappa_{r}}{\kappa_{t}}=\pm\dfrac{1}{f(r)}\sqrt{1-\dfrac{b^{2}}{r^{2}}f(r)},$ (41)
where the $\pm$ symbol indicates whether the photon is moving closer to or further away from the Schwarzschild-Hernquist black hole. Substituting these results into the expression for $\mathcal{R}$, the redshift factor becomes
$\mathcal{R}$ $=\left(u^{t}_{\text{emi}}+\dfrac{\kappa_{r}}{\kappa_{t}}u^{r}_{\text{emi}}\right)^{-1}$
$=\left[\dfrac{1}{f(r)}\pm\sqrt{\left(\dfrac{1}{f(r)}-1\right)\left(\dfrac{1}{f(r)}-\dfrac{b^{2}}{r^{2}}\right)}\right]^{-1}.$ (42)
Figure 11: The images of Schwarzschild-Hernquist black holes’ shaodow with infalling spherical accretion as observed by a distant observer.
Similarly, we maintain the assumption that the emission is monochromatic. Thus, for the infalling accrection, the observed intensity takes the form
$I_{\text{obs}}\propto\int_{\gamma}\dfrac{\mathcal{R}^{3}}{r^{2}}\dfrac{\kappa_{t}}{|\kappa_{r}|}\mathrm{d}r.$ (43)
The observed intensity profile $I_{\text{obs}}$ and the corresponding images for infalling accretion, computed using Eq.(43), are shown in Figs.10 and 11, respectively.
As shown in Fig.10, and similar to the static case, the observed intensity $I_{\text{obs}}$ for infalling accretion peaks near the critical impact parameter $b_{p}$ for all Hernquist DM parameter. The intensity increases for $b<b_{p}$, rising sharply just before reaching the peak at $b\sim b_{p}$. For $b<b_{p}$, the intensity gradually decreases with increasing $b$, consistent with the behavior in the static model. The observed intensity at $\rho_{c}M^{2}=0.4$ and $r_{s}/M=0.2$ is substantially brighter than at $\rho_{c}M^{2}=0.8$ and $r_{s}/M=0.4$, as shown in Figs.8 and 10. A direct comparison between the images in Fig.9 and Fig.11 reveals that the central region is significantly darker for infalling accretion. When the parameters are changed from $\rho_{c}M^{2}=0.4$ and $r_{s}/M=0.2$ to $\rho_{c}M^{2}=0.8$ and $r_{s}/M=0.4$, the peak intensity decreases by $18.36\%$ for static accretion and by $50.50\%$ for infalling accretion. As a direct result of the Doppler effect, the apparent intensity of radiation emitted by materials falling radially inward is reduced due to redshifting. Furthermore, in both the static and infalling models, the radius of the photon ring increases with increasing Hernquist DM parameters, consistent with the findings in Tab.1.
V Conclusion
In this work, we have conducted a thorough investigation of the optical appearance and shadows of a Schwarzschild black hole embedded within a Hernquist DM halo. Our findings demonstrate that the presence of a Hernquist DM halo substantially alters the observable characteristics of the black hole shadow across a variety of accretion scenarios. Employing ray-tracing techniques and numerical simulations, we have quantitatively analyzed three thin-disk accretion models (26)-(32), and two spherical accretion models Eqs.(37) and (43), revealing that the Hernquist DM parameters $\rho_{c}$ and $r_{s}$ have a significant impact on the photon sphere radius and the observed shadow properties. In particular, the photon sphere expands significantly as $\rho_{c}M^{2}$ and $r_{s}$ increase, as seen in Figs.6-11. Furthermore, we find that the core radius $r_{s}$ of the DM halo influences the growth of the photon sphere more significantly than the central density $\rho_{c}$. We use the Schwarzschild black hole with $b_{p}^{\text{Sch}}=3\sqrt{3}M$ as a reference caseLuminet (1979); Gralla et al. (2019). For the parameter set $\rho_{c}M^{2}=0.4$ and $r_{s}=0.2$, the shadow diameter increases relative to the Schwarzschild case by $\Delta\theta/\theta_{\text{Sch}}=\left(b_{p}^{\text{SH}}-b_{p}^{\text{Sch}}\right)/b_{p}^{\text{Sch}}=1.89\%$ Perlick and Tsupko (2022). Keeping $r_{s}=0.2$, the relative change rate $\Delta\theta/\theta_{\text{Sch}}$ rises to $3.78\%$ at $\rho_{c}M^{2}=0.8$. Notably, if we increase $r_{s}/M$ from $0.2$ to $0.4$ while keeping $\rho_{c}M^{2}=0.4$ fixed, the relative change climbs markedly to $14.40\%$. In extreme circumstances, the relative change rate $\Delta\theta/\theta_{\text{Sch}}$ reaches $29.15\%$ when $\rho_{c}M^{2}=0.8$ and $r_{s}=0.4$. However, as shown in Figs. 9 and 11, the peak observed intensity is significantly diminished, giving the overall image of being substantially dimmer. The gravitational influence of the Hernquist DM halo is responsible for these alterations, modifying the effective potential around the black hole and the photon trajectory. Our calculations indicate that while the presence of a Hernquist dark matter halo systematically increases the size of the shadow and the emission ring, the dynamical state of the accretion flow (static versus infalling) has a comparable, if not more significant, influence on the absolute level and distribution of brightness. This suggests that, when using EHT observations to constrain the central dark matter distribution, it is imperative to have a prior assumption or an independent constraint on the physical state of the accretion flow.
Importantly, our researches offer significant insights into potential observable features in real astrophysical systems like $\text{M87}^{*}$, which are typically described by the Kerr metric Collaboration (2019a, b, c, d, e, f, 2021a, 2021b, 2023), even though our model is based on a static, modified Schwarzschild background. The predicted variation in shadow size, which manifests as a positive correlation between the photon ring diameter and Hernquist DM parameters, offers a powerful diagnostic for probing the DM distribution around supermassive black holes. The current EHT observation for $\text{M87}^{*}$ is approximately $42\pm 3\,\mathrm{\mu as}$, whose relative uncertainty is about $7.14\%$ Collaboration (2019a), which is consistent with the Kerr vacuum solution. Our model provides an unambiguous prediction that may be tested by current or next-generation EHT measurements, since this stands in stark contrast to the $\sim 2\%$ to $\sim 30\%$ rise anticipated by Hernquist DM halos. Consequently, a future measurement of an anomalously large shadow could indicate the presence of a Hernquist DM halo, and our model could be used to interpret such a result. Conversely, the absence of significant shadow enlargement would place stringent constraints on the parameter space of viable Hernquist DM models.
Looking forward, our results highlight the potential of next-generation very-long-baseline interferometry (VLBI) Hobiger et al. (2009); Schuh and Böhm (2012); Niell et al. (2018); Nothnagel et al. (2018) arrays to refine DM models through precise shadow imaging. Furthermore, they reveal a novel, observable coupling mechanism between DM and black hole shadow phenomenology. Future work incorporating black hole spin like the Kerr metric Ferrer et al. (2017); Hou et al. (2018b) and more realistic, dynamic accretion flows will be essential to fully bridge the gap between theoretical predictions and observational evidence. Our results strongly support the use of black hole shadows as unique probes that can reveal the distribution and properties of DM halos, as well as elucidate new aspects of gravitational physics in strong-field regimes. This approach promises to illuminate the nature of DM in galactic centers in unprecedented ways.
Acknowledge
This work is partly supported by the Shanghai Key Laboratory of Astrophysics 18DZ2271600.